OpenStudy (ashy98):

x divided by -4 is 13. What is x???

OpenStudy (shamil98):

x divided BY -4 IS 13 is means equal divided by means x/n so x/-4 = 13 Multiply by the reciprocal of 1/-4 which is -4/1 to both sides x = 13 x -4 x = -52

OpenStudy (austinl):

\(\displaystyle \frac{x}{-4}=13\) To get rid of the negative 4 you need to multiply each side by -4. \(\displaystyle (\frac{-4}{1})(\frac{x}{-4})=13(-4)\) \(\displaystyle (\frac{\cancel{-4}}{1})(\frac{x}{\cancel{-4}})=13(-4)\) \(\displaystyle \frac{x}{1}=13\times-4=~?\)